need the trick for this problem: if the increase in the kinetic energy of a body is 22% then the increase in the momentum is nearly what %?

Grade:12th pass

1 Answers

Ashutosh nayak
7 years ago
Let 1st k.e. was100 then it increased by 22%so 2nd k.e ll b 122 which is almost equal to 121.as mass is constant and we know k.e.is 0.5 mv² so v2/v1 ll be (121/100)½=1.1.as momentum is directly proportional to vel when mass is constant..so %increase ll be ((v2-v1)/v1)*100 =((1.1-1)/1)*100 =10% so ur ans is 10%